# Covering rings by proper ideals

### v1

@@ Abstract.
Let $R$ be an associative (not necessarily commutative or unital) ring. A cover by left ideals of $R$ is a collection of proper left ideals whose set-theoretic union equals $R$. If such a cover exists, we define the left ideal covering number $\eta_{\ell}(R)$ to be the cardinality of a minimal cover, and we also define $R$ to be $\eta_{\ell}$-elementary if $\eta_{\ell}(R)<\eta_{\ell}(R/I)$ for every nonzero two-sided ideal $I$ of $R$. Similarly, we can define analogous notions for covers by right ideals $\eta_{r}(R)$ and two-sided ideals $\eta(R)$. In this paper, we classify all $\eta_{\ell}$-, $\eta_{r}$-, and $\eta$-elementary rings, and determine their corresponding ideal covering numbers. Since any ring $R$ that admits a finite cover by left ideals has an $\eta_{\ell}$-elementary residue ring $R/I$ such that $\eta_{\ell}(R)=\eta_{\ell}(R/I)$, this completely characterizes such rings and their ideal covering numbers, and likewise for rings coverable by right or two-sided ideals.
@@ 1. Introduction
The motivation for this problem comes from group theory. A group $G$ is said to be coverable if there exists a collection of proper subgroups of $G$ whose union is equal to $G$. It is an easy exercise to show that a group is never the union of two of its proper subgroups, and a classical result due to Scorza [Scorza_1926] shows that a group is the union of three proper subgroups if and only if it has a quotient isomorphic to the Klein four-group. Later, Cohn [Cohn_1994] began to systematically study the covering number $\sigma(G)$ of a group $G$, which is the minimum number of subgroups required to cover $G$. Cohn showed that there exists a group with covering number $p^{d}+1$ for every prime $p$ and integer $d\geqslant 1$, and Tomkinson [Tomkinson_1997] showed that there is no group with covering number seven. The covering numbers of groups have been studied extensively, and all integers $N\leqslant 129$ that are not the covering number of a group were determined in [Garonzi_Kappe_Swartz_2022], which also contains a survey of the related work for groups.
@@
This motivated inquiry into a broader question: What are the covering numbers of other algebraic structures? This has been studied for vector spaces [Clark_2012, Khare_2009], modules (over commutative rings) [Ghosh_2022, Khare_Tikaradze_2022], loops [Gagola_Kappe_2016], and semigroups [Donoven_Kappe_2023]. We refer the reader to the survey article [Kappe_2014] for more details.
@@
The related question for rings has also received significant attention. This was first initiated by Lucchini and Maróti [Lucchini_Maroti_2012], who determined all rings with covering number three. Building upon a series of earlier work [Cai_Werner_2019, Crestani_2012, Peruginelli_Werner_2018, Swartz_Werner_2021, Swartz_Werner_2024a], the recent article [Swartz_Werner_2024b] completely determined which integers can occur as covering numbers of rings. For a ring $R$, let $\sigma(R)$ be the minimum number of proper subrings of $R$ necessary to cover $R$ (if such a cover exists). The approach in [Swartz_Werner_2024b] is based on the observation that if $R$ has a two-sided ideal $I$ such that $R/I$ has a cover, then it can be lifted to a cover of $R$, so $\sigma(R)\leqslant\sigma(R/I)$. A ring $R$ is called $\sigma$-elementary if the inequality $\sigma(R)<\sigma(R/I)$ is strict for all $I\neq\{0\}$. Since any coverable ring is either $\sigma$-elementary, or has a proper $\sigma$-elementary residue ring with the same covering number, knowledge of $\sigma$-elementary rings allows us to characterize coverable rings and determine their covering numbers. A classification of all $\sigma$-elementary rings is given in [Swartz_Werner_2024b, Thm. 1.3].
@@
While many questions on the covering numbers of groups remain open, if one restricts to covers by normal subgroups there is an elegant result [Brodie_Chamberlain_Kappe_1988, Thm. 1] (see also [Bhargava_2002]): a group $G$ is the union of proper normal subgroups if and only if it has a quotient isomorphic to $C_{p}\times C_{p}$ (the direct product of two cyclic groups of prime order $p$); the corresponding covering number $\eta(G)$ is $p+1$ for the smallest such $p$. Our investigations in this paper are prompted by the ring-theoretic analogue, namely covers by proper ideals.
@@
Covers by two-sided ideals were first considered independently by O’Neill [O'Neill_1980] and Parmenter [Parmenter_1994]; in each paper, the authors provide characterizations of rings with covers by two-sided ideals. In contrast, covers by left (resp. right) ideals were, to the best of our knowledge, first studied in the preprint [Chen_2025], which suggested that
@@
From Theorem 1.2, we recover [O'Neill_1980, Thm. 1] and [Parmenter_1994, Thm. 2]: $R$ is coverable by finitely many two-sided ideals if and only if $R$ has a finite homomorphic image $S$ such that $S^{2}=\{0\}$ and $(S,+)$ is a nontrivial noncyclic group. We remark that Theorem 1.2 can be proved independently by appealing to [O'Neill_1980, Thm. 1] or [Parmenter_1994, Thm. 2]—which reduces the problem to covering a noncyclic abelian group—and applying a result of Cohn [Cohn_1994, Thm. 2]. We include our proof of Theorem 1.2 for completeness.
@@
This generalizes the known results on covers of modules over commutative rings. In [Khare_Tikaradze_2022, Thm. 2.2], it is shown that if $R$ is a commutative unital ring and $M$ is an $R$-module that admits a finite cover by proper $R$-submodules, then the covering number of $M$ is $q+1$ for some prime power $q$. Our work shows that the determination of covering numbers of modules is still an open problem. It would be interesting to know if there are other integers that are covering numbers of modules.
@@
To prove Theorem 1.3(1), we show that $S\cong M_{n}(\mathbb{F}_{q})$ and $J\cong\mathbb{F}_{q}^{n}$ for some prime power $q$ and integer $n\geqslant 1$. In Section 4, we show that when $S$ and $J$ have these forms, the ring $R$ in (1.1) is $\eta_{\ell}$-elementary and $\eta_{\ell}(R)=\frac{q^{n+1}-1}{q-1}$. In Section 5, we prove that these are the only noncommutative $\eta_{\ell}$-elementary rings.
@@ 2. Preliminaries
If $R$ is finite, then the structure of $R^{\prime}$ is determined by the Wedderburn-Malcev Theorem (sometimes also called the Wedderburn Principal Theorem); see e.g. [Pierce_1982, Sec. 11.6, Cor. p. 211] or [McDonald_1974, Thm. VIII.28]. To describe the theorem, we need to define the Jacobson radical of a ring. There are a number of common equivalent definitions for the Jacobson radical of a unital ring. For a nonunital ring, the definition is more obscure. We follow the approach taken in [Lam_1991, Sec. 4].
If we define the operation $\circ$ on a (not necessarily unital) ring $T$ by $a\circ b\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}a+b-ab$, then $(T,\circ)$ is a monoid with 0 as the identity element. An element of $T$ is called left (resp. right) quasi-regular if it has a left (resp. right) inverse in the monoid $(T,\circ)$. A subset $I\subseteq T$ is called left (resp. right) quasi-regular if every element of $I$ is left (resp. right) quasi-regular. As in [Lam_1991, Sec. 4, Exercise 4], we have the following definition.
@@
(Nakayama’s Lemma) [Lam_1991, Lem. 4.22] Let $T$ be a unital ring and $I$ be a left ideal of $T$. Then, $I\subseteq\mathscr{J}(T)$ if and only if for any finitely generated left $T$-module $M$ and $N$ a submodule of $M$, $M=N+IM$ implies that $M=N$.
@@
The following proof is adapted from and analogous to that of [rschwieb_2016]. First, since $\{0\}\times R$ is a maximal ideal of $R^{\prime}$ by Lemma 2.1, $\mathscr{J}(R^{\prime})\subseteq\{0\}\times R$.
@@
so $R/J$ is isomorphic to an ideal of the semisimple ring $S^{\prime}$. By [McDonald_1974, Thm. VIII.5], $R/J$ is either $\{0\}$ or semisimple. Since $S^{\prime}\cap(\{0\}\times J)=\{0_{R^{\prime}}\}$, there exists an ideal $\{0\}\times S$ of $S^{\prime}$ such that
@@ 3. Basic properties and reduction theorems
We begin with a variation on [Werner_2015, Thm. 2.2], which gives a sufficient condition for the covering number of a direct product of rings to be equal to the covering number of one of the direct factors.
@@
(2) This is identical to the proof of [Werner_2015, Thm. 2.2] with each use of “maximal subring” replaced with “maximal *-ideal”.
@@
(1) Suppose that $\eta_{*}(R)<\infty$ and $R=\bigcup_{i=1}^{\eta_{*}(R)}L_{i}$ is a minimal cover by *-ideals. By [Neumann_1954, Lem. 4.1, 4.4], each $(L_{i},+)$ has finite index in $(R,+)$, and by [Lewin_1967, Lem. 1], their intersection contains a two-sided ideal $I$ of finite index in $R$, so $R/I$ is finite. Since the images of $L_{i}$ mod $I$ form a *-ideal cover of $R/I$, and conversely any *-ideal cover of $R/I$ can be lifted to a *-ideal cover of $R$, we have $\eta_{*}(R)=\eta_{*}(R/I)$.
(2) Assume that $R$ is $\eta_{*}$-elementary. By (1), $R$ must be finite, because the condition $\eta_{*}(R)=\eta_{*}(R/I)$ forces $I=\{0\}$. It is well known that any finite (not necessarily unital) ring $R$ is isomorphic to a direct product of rings of prime power order, so $R=\prod_{i=1}^{t}R_{i}$ where $|R_{i}|=p_{i}^{d_{i}}$ for distinct primes $p_{i}$. It follows from Lemma 3.1(3) that $R$ has prime power order and hence has characteristic $p^{d}$ for some prime $p$ and integer $d\geqslant 1$. We now argue as in the proof of [Swartz_Werner_2021, Lem. 3.2(2), Prop. 3.3]: $pR$ is a two-sided ideal of $R$, and, if $M$ is a maximal *-ideal of $R$, $M+pR$ is also a *-ideal. Since $M\subseteq M+pR\subseteq R$, by the maximality of $M$, either $M+pR=M$ or $M+pR=R$. However, by [Swartz_Werner_2021, Lem. 3.2(1)] (which holds for nonunital rings), $M+pR=R$ implies $M=R$, contradicting our assumptions. Thus, $pR\subseteq M$ for all maximal *-ideals $M$. In particular, if $L_{1},\dots,L_{m}$ is a minimal cover of $R$ by *-ideals, without loss of generality we may assume each $L_{i}$ is a maximal *-ideal, and hence $pR\subseteq L_{i}$ for each $i$. By Lemma 3.2, $pR=\{0\}$, and so $R$ has characteristic $p$. ∎
@@
By Lemma 3.3, $R$ has characteristic $p$. Suppose $\mathscr{J}(R)=\{0\}$. Then, $\mathscr{J}(R^{\prime})=\{0\}$ by Lemma 2.6, and so $R^{\prime}$ is semisimple by Theorem 2.4. By Lemma 2.1, $\{0\}\times R$ is a two-sided ideal of $R^{\prime}$, so $R$ is a direct sum of simple direct factors of $R^{\prime}$; see e.g. [McDonald_1974, Thm. VIII.5]. However, this implies that $R$ contains a multiplicative identity, which is a contradiction to $R$ having a cover by *-ideals. ∎
@@
We claim that $S=\{0\}$. Suppose not, and let $e=1_{S}$. If $(J,+)=\langle x\rangle$ is cyclic, then $R$ is not coverable. Indeed, by (3.1) we have $e(e+x)=e^{2}=e$, so any ideal of $R$ containing $e+x$ contains both $e$ and $x$, and hence must equal $R$. So, $(J,+)$ is not cyclic, and hence $J$ is coverable by two-sided ideals. Since ideal covers of $J$ correspond to subgroup covers of $(J,+)$, we have $\eta(J)=p+1$ by [Cohn_1994, Thm. 2]. But then, $\eta(R)<\eta(R/S)=\eta(J)=p+1$ because $R$ is $\eta$-elementary. This is impossible, because an ideal cover of $R$ induces an additive cover of $(R,+)$, and $(R,+)$ has covering number $p+1$ by [Cohn_1994, Thm. 2]. We conclude that $S=\{0\}$ and $R=J$. As a consequence, we have $R^{2}=\{0\}$.
@@
The rings described Theorem 1.3(1) have the form specified by Proposition 3.8(2). To prove Theorem 1.3(1), it remains to show two things. First, that there exist $\eta_{\ell}$-elementary rings of the form given in Proposition 3.8(2). Second, when $R=S\oplus J$ is $\eta_{\ell}$-elementary, we must prove that $S$ is a simple ring and $J$ is a simple left $S$-module. These claims are dealt with in Sections 4 and 5.

### v2

@@ Abstract.
A cover by left ideals of an associative (not necessarily commutative or unital) ring $R$ is a collection of proper left ideals whose set-theoretic union equals $R$. If such a cover exists, then $\eta_{\ell}(R)$ is the cardinality of a minimal cover, and $R$ is $\eta_{\ell}$-elementary if $\eta_{\ell}(R)<\eta_{\ell}(R/I)$ for every nonzero two-sided ideal $I$ of $R$. We classify all $\eta_{\ell}$-elementary rings, and determine their covering numbers. Covers by right or two-sided ideals are also studied. This completely characterizes rings admitting finite covers by ideals. Our results generalize to finite covers of modules by submodules, and we determine all possible covering numbers.
@@ 1. Introduction
The motivation for this problem comes from group theory. A group $G$ is said to be coverable if there exists a collection of proper subgroups of $G$ whose union is equal to $G$. It is an easy exercise to show that a group is never the union of two of its proper subgroups, and a classical result due to Scorza [25] shows that a group is the union of three proper subgroups if and only if it has a quotient isomorphic to the Klein four-group. Later, Cohn [6] began to systematically study the covering number $\sigma(G)$ of a group $G$, which is the minimum number of subgroups required to cover $G$. Cohn showed that there exists a group with covering number $p^{d}+1$ for every prime $p$ and integer $d\geqslant 1$, and Tomkinson [29] showed that there is no group with covering number seven. The covering numbers of groups have been studied extensively, and all integers $N\leqslant 129$ that are not the covering number of a group were determined in [10], which also contains a survey of the related work for groups.
@@
This motivated inquiry into a broader question: What are the covering numbers of other algebraic structures? This has been studied for vector spaces [5, 14], modules (over commutative rings) [11, 13], loops [9], and semigroups [8]. We refer the reader to the survey article [12] for more details.
@@
The related question for rings has also received significant attention. This was first initiated by Lucchini and Maróti [17], who determined all rings with covering number three. Building upon a series of earlier work [3, 7, 22, 26, 27], the recent article [28] completely determined which integers can occur as covering numbers of rings. For a ring $R$, let $\sigma(R)$ be the minimum number of proper subrings of $R$ necessary to cover $R$ (if such a cover exists). The approach in [28] is based on the observation that if $R$ has a two-sided ideal $I$ such that $R/I$ has a cover, then it can be lifted to a cover of $R$, so $\sigma(R)\leqslant\sigma(R/I)$. A ring $R$ is called $\sigma$-elementary if the inequality $\sigma(R)<\sigma(R/I)$ is strict for all $I\neq\{0\}$. Since any coverable ring is either $\sigma$-elementary, or has a proper $\sigma$-elementary residue ring with the same covering number, knowledge of $\sigma$-elementary rings allows us to characterize coverable rings and determine their covering numbers. A classification of all $\sigma$-elementary rings is given in [28, Thm. 1.3].
@@
While many questions on the covering numbers of groups remain open, if one restricts to covers by normal subgroups there is an elegant result [2, Thm. 1] (see also [1]): a group $G$ is the union of proper normal subgroups if and only if it has a quotient isomorphic to $C_{p}\times C_{p}$ (the direct product of two cyclic groups of prime order $p$); the corresponding covering number $\eta(G)$ is $p+1$ for the smallest such $p$. Our investigations in this paper are prompted by the ring-theoretic analogue, namely covers by proper ideals.
@@
Covers by two-sided ideals were first considered independently by O’Neill [20] and Parmenter [21]; in each paper, the authors provide characterizations of rings with covers by two-sided ideals. In contrast, covers by left (resp. right) ideals were, to the best of our knowledge, first studied in the preprint [4], which suggested that
@@
From Theorem 1.2, we recover [20, Thm. 1] and [21, Thm. 2]: $R$ is coverable by finitely many two-sided ideals if and only if $R$ has a finite homomorphic image $S$ such that $S^{2}=\{0\}$ and $(S,+)$ is a nontrivial noncyclic group. We remark that Theorem 1.2 can be proved independently by appealing to [20, Thm. 1] or [21, Thm. 2]—which reduces the problem to covering a noncyclic abelian group—and applying a result of Cohn [6, Thm. 2]. We include our proof of Theorem 1.2 for completeness.
@@
This generalizes the known results on covers of modules over commutative rings. In [13, Thm. 2.2], it is shown that if $R$ is a commutative unital ring and $M$ is an $R$-module that admits a finite cover by proper $R$-submodules, then the covering number of $M$ is $q+1$ for some prime power $q$. We are able to generalize these results to modules over noncommutative rings, and prove that the covering numbers obtained in Corollary 1.4 are the only possible finite covering numbers for modules.
Theorem 1.5.
Let $R$ be a unital ring and let $M$ be a left (resp. right) $R$-module that admits a finite cover by proper left (resp. right) $R$-submodules. Then, there exists some prime power $q$ and integer $n\geqslant 1$ such that the covering number of $M$ by left (resp. right) $R$-modules is $\tfrac{q^{n+1}-1}{q-1}$.
@@
To prove Theorem 1.3(1), we show that $S\cong M_{n}(\mathbb{F}_{q})$ and $J\cong\mathbb{F}_{q}^{n}$ for some prime power $q$ and integer $n\geqslant 1$. In Section 4, we show that when $S$ and $J$ have these forms, the ring $R$ in (1.1) is $\eta_{\ell}$-elementary and $\eta_{\ell}(R)=\frac{q^{n+1}-1}{q-1}$. In Section 5, we prove that these are the only noncommutative $\eta_{\ell}$-elementary rings. In Section 6, we close the paper with the proof of Theorem 1.5.
@@ 2. Preliminaries
If $R$ is finite, then the structure of $R^{\prime}$ is determined by the Wedderburn-Malcev Theorem (sometimes also called the Wedderburn Principal Theorem); see e.g. [23, Sec. 11.6, Cor. p. 211] or [18, Thm. VIII.28]. To describe the theorem, we need to define the Jacobson radical of a ring. There are a number of common equivalent definitions for the Jacobson radical of a unital ring. For a nonunital ring, the definition is more obscure. We follow the approach taken in [15, Sec. 4].
If we define the operation $\circ$ on a (not necessarily unital) ring $T$ by $a\circ b\mathrel{\mathchoice{\vbox{\hbox{$\displaystyle:$}}}{\vbox{\hbox{$\textstyle:$}}}{\vbox{\hbox{$\scriptstyle:$}}}{\vbox{\hbox{$\scriptscriptstyle:$}}}{=}}a+b-ab$, then $(T,\circ)$ is a monoid with 0 as the identity element. An element of $T$ is called left (resp. right) quasi-regular if it has a left (resp. right) inverse in the monoid $(T,\circ)$. A subset $I\subseteq T$ is called left (resp. right) quasi-regular if every element of $I$ is left (resp. right) quasi-regular. As in [15, Sec. 4, Exercise 4], we have the following definition.
@@
(Nakayama’s Lemma) [15, Lem. 4.22] Let $T$ be a unital ring and $I$ be a left ideal of $T$. Then, $I\subseteq\mathscr{J}(T)$ if and only if for any finitely generated left $T$-module $M$ and $N$ a submodule of $M$, $M=N+IM$ implies that $M=N$.
@@
The following proof is adapted from and analogous to that of [24]. First, since $\{0\}\times R$ is a maximal ideal of $R^{\prime}$ by Lemma 2.1, $\mathscr{J}(R^{\prime})\subseteq\{0\}\times R$.
@@
so $R/J$ is isomorphic to an ideal of the semisimple ring $S^{\prime}$. By [18, Thm. VIII.5], $R/J$ is either $\{0\}$ or semisimple. Since $S^{\prime}\cap(\{0\}\times J)=\{0_{R^{\prime}}\}$, there exists an ideal $\{0\}\times S$ of $S^{\prime}$ such that
@@ 3. Basic properties and reduction theorems
We begin with a variation on [30, Thm. 2.2], which gives a sufficient condition for the covering number of a direct product of rings to be equal to the covering number of one of the direct factors.
@@
(2) This is identical to the proof of [30, Thm. 2.2] with each use of “maximal subring” replaced with “maximal *-ideal”.
@@
(1) Suppose that $\eta_{*}(R)<\infty$ and $R=\bigcup_{i=1}^{\eta_{*}(R)}L_{i}$ is a minimal cover by *-ideals. By [19, Lem. 4.1, 4.4], each $(L_{i},+)$ has finite index in $(R,+)$, and by [16, Lem. 1], their intersection contains a two-sided ideal $I$ of finite index in $R$, so $R/I$ is finite. Since the images of $L_{i}$ mod $I$ form a *-ideal cover of $R/I$, and conversely any *-ideal cover of $R/I$ can be lifted to a *-ideal cover of $R$, we have $\eta_{*}(R)=\eta_{*}(R/I)$.
(2) Assume that $R$ is $\eta_{*}$-elementary. By (1), $R$ must be finite, because the condition $\eta_{*}(R)=\eta_{*}(R/I)$ forces $I=\{0\}$. It is well known that any finite (not necessarily unital) ring $R$ is isomorphic to a direct product of rings of prime power order, so $R=\prod_{i=1}^{t}R_{i}$ where $|R_{i}|=p_{i}^{d_{i}}$ for distinct primes $p_{i}$. It follows from Lemma 3.1(3) that $R$ has prime power order and hence has characteristic $p^{d}$ for some prime $p$ and integer $d\geqslant 1$. We now argue as in the proof of [26, Lem. 3.2(2), Prop. 3.3]: $pR$ is a two-sided ideal of $R$, and, if $M$ is a maximal *-ideal of $R$, $M+pR$ is also a *-ideal. Since $M\subseteq M+pR\subseteq R$, by the maximality of $M$, either $M+pR=M$ or $M+pR=R$. However, by [26, Lem. 3.2(1)] (which holds for nonunital rings), $M+pR=R$ implies $M=R$, contradicting our assumptions. Thus, $pR\subseteq M$ for all maximal *-ideals $M$. In particular, if $L_{1},\dots,L_{m}$ is a minimal cover of $R$ by *-ideals, without loss of generality we may assume each $L_{i}$ is a maximal *-ideal, and hence $pR\subseteq L_{i}$ for each $i$. By Lemma 3.2, $pR=\{0\}$, and so $R$ has characteristic $p$. ∎
@@
By Lemma 3.3, $R$ has characteristic $p$. Suppose $\mathscr{J}(R)=\{0\}$. Then, $\mathscr{J}(R^{\prime})=\{0\}$ by Lemma 2.6, and so $R^{\prime}$ is semisimple by Theorem 2.4. By Lemma 2.1, $\{0\}\times R$ is a two-sided ideal of $R^{\prime}$, so $R$ is a direct sum of simple direct factors of $R^{\prime}$; see e.g. [18, Thm. VIII.5]. However, this implies that $R$ contains a multiplicative identity, which is a contradiction to $R$ having a cover by *-ideals. ∎
@@
We claim that $S=\{0\}$. Suppose not, and let $e=1_{S}$. If $(J,+)=\langle x\rangle$ is cyclic, then $R$ is not coverable. Indeed, by (3.1) we have $e(e+x)=e^{2}=e$, so any ideal of $R$ containing $e+x$ contains both $e$ and $x$, and hence must equal $R$. So, $(J,+)$ is not cyclic, and hence $J$ is coverable by two-sided ideals. Since ideal covers of $J$ correspond to subgroup covers of $(J,+)$, we have $\eta(J)=p+1$ by [6, Thm. 2]. But then, $\eta(R)<\eta(R/S)=\eta(J)=p+1$ because $R$ is $\eta$-elementary. This is impossible, because an ideal cover of $R$ induces an additive cover of $(R,+)$, and $(R,+)$ has covering number $p+1$ by [6, Thm. 2]. We conclude that $S=\{0\}$ and $R=J$. As a consequence, we have $R^{2}=\{0\}$.
@@
The rings described Theorem 1.3(1) have the form specified by Proposition 3.8. To prove Theorem 1.3(1), it remains to show two things. First, that there exist $\eta_{\ell}$-elementary rings of the form given in Proposition 3.8. Second, when $R=S\oplus J$ is $\eta_{\ell}$-elementary, we must prove that $S$ is a simple ring and $J$ is a simple left $S$-module. These claims are dealt with in Sections 4 and 5.